OpenStudy (anonymous):

what is the eqaution of the line. x goes truh 7 and y goes truh -4

OpenStudy (anonymous):

????????

OpenStudy (anonymous):

A. –4x + 7y = 28 B. 4x + 7y = –28 C. –4x – 7y = –28 D. 4x – 7y = 28

OpenStudy (anonymous):

will give meadl!!!!!!!!!!!!!

OpenStudy (anonymous):

*medal*!!!!!!!!!!!!!!!!!!!!!

OpenStudy (eric_d):

confirm it's not A hang on

OpenStudy (eric_d):

not C.

OpenStudy (eric_d):

D.